Shift report 20010820 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 20 2001 - 00:05:21 EDT

  • Next message: Apache: "Shift report 20010820 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Aug 20 00:05:20 2001, Shift: 16:00-24:00
    Supervisor: JJG, with: AM
    Log book#7 pages 137 - 137
    
    Shift Summary:
    ------------------------------------------------------------
    1. Collisions establisehd at around 50 evts/sec from 
      16.5 to 19.45 when beam aborted due to magnet quench.
      Since then no beam. Injection problems at AGS. Current RHIC
      prediction for startup not clear.
    2.Run 4295 contains about 361K trigger 1's. 
      Runs 4294 and 4296 are ped. runs.
    3.We didn't fix any detectors, nor destroyed any, but 
      forgot to turn T1 anode up (jjg screwed up, but its a 
      bad idea to have the standby voltage show up orange 
      and not red). 
    4.Should request access on monday to install T1 pulser and 
      work on T3 module 2.
    5.We now should have about 1M events at 4 deg pol A.
      and around 840k trig 1's at pol B (although 361 K with 
      T1).  
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 20 2001 - 00:06:15 EDT